# Mapping bipartite networks into multidimensional hyperbolic spaces

### v1

@@ IV.2 B-Mercator in details
The inference of hidden degrees for type A and B nodes, and the inverse temperature $\beta_{b}$ is implemented as an iterative process. We begin with the initial guess for the parameter $\beta_{b}\in(D,2D)$, where $D$ is an embedding dimension, and initialize the hidden degrees as the observed degrees in the original network for nodes A ($\{k_{A,i},i,\dots,N_{A}\}$) and nodes B ($\{k_{B,i},i,\dots,N_{B}\}$). The aim of the estimation is to modify the hidden degrees in order to ensure that the expected degree of each node within the model aligns with the degree observed in the original network. After the hidden degrees for both nodes A and B are computed, the synthetic graph from bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ is constructed and the bipartite clustering coefficient is calculated. If the computed bipartite clustering coefficient deviates from that of the original network, $\bar{c}_{b}$, the value of $\beta_{b}$ is adjusted. Then, the process is repeated using the current estimation of hidden degrees until a predetermined precision is reached.
@@
Since the biadjacency matrix of a bipartite graph $\mathbf{A}$ is not symmetric, to apply the Laplacian Eigenmaps, we transform it to the adjacency matrix as
@@
Similarly to Jankowski et al. (2023), the expected angular distance between nodes $u$ (of type A) and $v$ (of type B) in the bipartite-$\mathbb{S}^{D}$ model, conditioned to the fact that they are connnected, can be computed as
[…]
Additionally, for $D=1$, we keep the ordering inferred by LE and distribute nodes of type A and B evenly on the circle.
@@
The nodes’ coordinates in the similarity space inferred using LE are adjusted by Maximum Likelihood Estimation (MLE) to optimize the probability that the bipartite-$\mathbb{S}^{D}$ model generates the observed network. We define an order of nodes sorted by their degree for A and B type nodes separately. Fixing the positions of a subset of B type nodes, we find new optimal coordinates for a subset of type A nodes. First, we compute the mean coordinates of type A node $u$’s neighbors.
@@
where the sum goes of all neighbors of node $u$, i.e., nodes of type B. Later, the new positions around $\mathbf{x}_{u}$ are proposed using a multivariate normal distribution. Finally, we select the most likely candidate position based on the local log-likelihood
@@
After iterating over a subset of nodes of type A, we apply a similar approach to nodes of type B, and repeat the process until all node positions are adjusted.
@@
Lastly, we adjust the hidden degree to compensate deviations from $\bar{k}_{A}(\kappa_{u})=\kappa_{u}$ and $\bar{k}_{B}(\kappa_{v})=\kappa_{v}$, which might have been introduced during the estimation of the nodes’ coordinates in the similarity space.
IV.3 Real bipartite datasets
The Unicodelang dataset was downloaded from the Unicode CLDR Project GitHub repository Consortium (2024). The dataset [^1] https://github.com/unicode-org/cldr/blob/main/common/supplemental/supplementalData.xml contains information about the languages spoken in a given territory. We preprocessed this dataset and matched the country codes to their geographical regions [^2] https://github.com/lukes/ISO-3166-Countries-with-Regional-Codes. In total, the bipartite graph contains 246 countries, 717 languages, and 1487 edges.
@@ Acknowledgments
We acknowledge support from: Grant TED2021-129791B-I00 funded by MCIN/AEI/10.13039/501100011033 and the “European Union NextGenerationEU/PRTR”; Grant PID2019-106290GB-C22 funded by MCIN/AEI/10.13039/501100011033; Generalitat de Catalunya grant number 2021SGR00856. R. J. acknowledges support from the fellowship FI-SDUR funded by Generalitat de Catalunya. M. B. acknowledges the ICREA Academia award, funded by the Generalitat de Catalunya.
V Data availability
@@
VI Code availability
The open-source code for B-Mercator, along with the code to reproduce the figures, will available on GitHub at https://github.com/networkgeometry/b-mercator upon publication.

### v2

@@ III.2 B-Mercator in detail
The inference of hidden degrees for type-A and -B nodes, and the inverse temperature $\beta_{b}$ is implemented as an iterative process. We begin with the initial guess for the parameter $\beta_{b}\in(D,2D)$ where $D$ is the embedding dimension, and initialize the hidden degrees as the observed degrees in the original network for type-A ($k_{A}=\{k_{A,i}\}_{i=1}^{N_{A}}$) and type-B ($k_{B}=\{k_{B,j}\}_{j=1}^{N_{B}}$) nodes. The aim of the estimation is to modify the hidden degrees in order to ensure that the expected degree of each node within the model aligns with the degree observed in the original network. After the hidden degrees for both nodes A and B are computed, the synthetic graph from the bipartite-$\mathbb{S}^{D}/\mathbb{H}^{D+1}$ is constructed and the bipartite clustering coefficient is calculated. If the computed bipartite clustering coefficient deviates from that of the original network, $\bar{c}_{b}$, the value of $\beta_{b}$ is adjusted. Then, the process is repeated using the current estimation of hidden degrees until a predetermined precision is reached.
@@
Since the biadjacency matrix of a bipartite graph $\mathbf{A}$ is not symmetric, to apply the Laplacian Eigenmaps Belkin and Niyogi (2003), we transform it to the adjacency matrix as
@@
Similarly to Jankowski et al. (2023), the expected angular distance between nodes $u$ (of type-A) and $v$ (of type-B) in the bipartite-$\mathbb{S}^{D}$ model, conditioned to the fact that they are connected, can be computed as
$\langle\Delta\theta_{uv}\rangle$ $=\int\limits_{0}^{\pi}\Delta\theta_{uv}\rho(\Delta\theta_{uv}|a_{uv}=1)\,d\Delta\theta_{uv}$ (7)
Additionally, for $D=1$, we keep the ordering inferred by LE and distribute type-A and type-B nodes evenly on the circle.
@@
The nodes’ coordinates in the similarity space inferred using LE are adjusted by Maximum Likelihood Estimation (MLE) to optimize the probability that the bipartite-$\mathbb{S}^{D}$ model generates the observed network. We define an order of nodes sorted by their degree for A and B type nodes separately. Fixing the positions of a subset of B type nodes, we find new optimal coordinates for a subset of type-A nodes. First, we compute the mean coordinates of type-A node $u$’s neighbors.
@@
where the sum goes over all neighbors of node $u$, i.e., type-B nodes. Later, the new positions around $\mathbf{x}_{u}$ are proposed using a multivariate normal distribution. Finally, we select the most likely candidate position based on the local log-likelihood
@@
After iterating over a subset of type-A nodes, we apply a similar approach to type-B nodes, and repeat the process until all node positions are adjusted.
@@
Lastly, we adjust the hidden degrees to compensate deviations from $\bar{k}_{A}(\kappa_{u})=\kappa_{u}$ and $\bar{k}_{B}(\kappa_{v})=\kappa_{v}$, which might have been introduced during the estimation of the nodes’ coordinates in the similarity space.
III.3 Real bipartite datasets
The Unicodelang dataset was downloaded from the Unicode CLDR Project GitHub repository Unicode Consortium (2024). The dataset The Unicode Consortium (2025) contains information about the languages spoken in a given territory. We preprocessed this dataset and matched the country codes to their geographical regions Duncalfe (2025). In total, the bipartite graph contains 246 countries, 717 languages, and 1487 edges.
@@ Acknowledgments
We acknowledge support from: Grant TED2021-129791B-I00 funded by MCIN/AEI/10.13039/501100011033 and the “European Union NextGenerationEU/PRTR”; Grant PID2022-137505NB-C22 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF/EU”. R. J. acknowledges support from the fellowship FI-SDUR funded by Generalitat de Catalunya. M. B. acknowledges the ICREA Academia award, funded by the Generalitat de Catalunya.
Author Contributions
M. B., M. A. S., R. J., and R. A. designed research. R. J., and R. A. performed research. M. B., M. A. S., R. J. wrote the paper. All authors discussed results, reviewed the manuscript, and approved the final version.
Competing interests
All authors declare no competing interests.
Data availability
@@
Code availability
The open-source code for B-Mercator, along with the code to reproduce the figures, is available on GitHub networkgeometry (2025).